Two angles are complementary if the sum of their measures is 90°. The measure of the first angle is 5° less than 4 times the second angle. Find the measure of the FIRST angle.
Answer by rapaljer(4667)

Let x = second angle
4x-5 = first angle

x + 4x-5 = 90
5x-5 = 90
5x=95
x=19 Second angle

4x-5 = First angle
4(19) - 5=76-5 = 71 = First angle

Check:
19+71=90

R^2